mhchen:

Right Riemann sum for f(x) on [0.3], n>0 n = number of bars f(x) = (x^3)/9 Calculate the Riemann for n.

mhchen:

\[\sum_{k=1}^{n}f(x)\Delta x\] I need to calculate this formula by getting rid of the summation. I know \[\Delta x = \frac{(3-0)}{n} = \frac{3}{n}\] and \[ f(x) = \frac{x^3}{9} \] So I plug them in to get \[\sum_{k=1}^{n} \frac{x^3}{9} \frac{3}{n}\]

Ultrilliam:

@563blackghost @Vocaloid

mhchen:

\[\sum_{k=1}^{n}Ck^3 = C\frac{n^2(n+1)^2}{4}\] They gave us this formula to work with as well. Oh sorry I should also mention that x represents the grid points so it should be x subbed k. \[x_{k}\]

mhchen:



mhchen:

Oof nvm I solved it myself. I see what I did wrong.